OpenStudy (anonymous):

activitity title> constant velocity speed exercises Motion a line with constant velocity. 1.if it takes 2 hours for a bus to travel 100 km along a straight highway what is its speed? 2.if a van travel 72 km on a straight expressway for 240 minutes,what is its speed? 3.a jeepney or a car at a constant speed of 40km/hr. how many minutes will it take for the jeepney to travel a distance 0f 80 km?

OpenStudy (anonymous):

1)distance=100km, time=2 hours speed=distance/time=100/2=50km/hr 2)distance=72 km, time=240mts=240/60hrs=4hrs speed=distance/time=72/4=18km/hr 3)speed =40km/hr, distance =80km speed=distance/time time=distance/speed=80/40=2hrs